You receive a $165,000 inheritance from your parents and invest it in a savings plan earning 6 5/8% interest compounded monthly. After 3 years you decide to deposit an additional $12,353.89 in the account and let the money sit for 10 more years. What is the future value of this investment?
Answer by Theo(13342)

your annual interest rate is 6 and 5/8 percent.

divide that by 12 to get a monthly interest rate of 53/96 percent per month.

divide that by 100 and the interest rate is 53/9600 per month.

your initial investment is 165,000 which sits there for 3 + 10 = 13 years.

your additional investment is 12,353.89 which sits there for 10 years.

13 years * 12 = 156 months.

10 years * 12 = 120 months.

the future value of the 165,000 is equal to 165,000 * (1 + 53/9600) ^ 156 = 389,485.5552.

the future value of the 12,353.89 is equal to 12,353.89 * (1 + 53/9600) ^ 120 = 23,918.42112

your total future value is the sum of these which is equal to 413,403.9763.